### v2

# CARScenes: Semantic VLM Dataset for Safe Autonomous Driving

Abstract
CARScenes is a frame-level dataset for autonomous driving that enables the training and evaluation of vision–language models (VLMs) for interpretable, scene-level understanding. We annotate 5,192 images drawn from Argoverse1, Cityscapes, KITTI, and nuScenes using a 28-key category/sub-category knowledge base covering environment, road geometry, background-vehicle behavior, ego-vehicle behavior, vulnerable road users, sensor states, and a discrete severity scale (1–10), totaling 350+ leaf attributes. Labels are produced by a GPT-4o–assisted vision–language pipeline with human-in-the-loop verification; we release the exact prompts, post-processing rules, and per-field baseline model performance. CARScenes also provides attribute co-occurrence graphs and JSONL records that support semantic retrieval, dataset triage, and risk-aware scenario mining across sources. To calibrate task difficulty, we include reproducible, non-benchmark baselines—notably a LoRA-tuned Qwen2-VL-2B with deterministic decoding—evaluated via scalar accuracy, micro-averaged F1 for list attributes, and severity MAE/RMSE on a fixed validation split. We publicly release the annotation, and analysis scripts, including graph construction and evaluation scripts, to enable explainable, data-centric workflows for future intelligent vehicles.
I Introduction
Recent progress in autonomous driving has been propelled by large-scale datasets that capture diverse environments, multimodal sensors, and richly annotated objects and lanes. Foundational benchmarks such as KITTI [RefWorks:geiger2013vision], Cityscapes [RefWorks:cordts2016cityscapes], nuScenes [RefWorks:caesar2020nuscenes], Argoverse [RefWorks:chang2019argoverse], Waymo Open [RefWorks:sun2020scalability], and BDD100K [RefWorks:yu2020bdd100k] have enabled reproducible research across core tasks—2D/3D detection, segmentation, tracking, motion forecasting, and trajectory prediction. Yet, most emphasize low-level perception and spatial reasoning, providing bounding boxes, masks, or trajectories, with limited coverage of high-level scene semantics (e.g., road context, environmental conditions, agent behaviors) that shape human driving decisions.
More recently, as the field increasingly explores vision–language models (VLMs) and language-conditioned interfaces for autonomous driving—spanning language grounding and instruction following [RefWorks:deruyttere2019talk2car] to multi-step reasoning and QA about scenes [RefWorks:sima2024drivelm]—there is a growing need for structured, interpretable scene-level annotations that can both train and evaluate VLMs. General-purpose VLMs (e.g., LLaVA [RefWorks:liu2023visual], Qwen-VL [RefWorks:wang2024qwen2vl]) demonstrate strong multimodal priors, but lack domain-specific supervision that encodes the semantics and safety cues unique to driving. As a result, we lack per-image, schema-aligned scene-level annotations that are suitable for training and evaluating vision-language models beyond free-text QA or narrow task labels.
We introduce CARScenes, a frame-level dataset designed for training and evaluating VLMs on interpretable scene understanding. CARScenes consolidates 5,192 images from Argoverse1, Cityscapes, KITTI, and nuScenes under a 28-key category/sub-category knowledge base ( 350+ leaf attributes) that covers environment, road geometry, background-vehicle behavior, ego-vehicle behavior, vulnerable road users, sensor states, and a discrete severity scale (1–10). Labels are produced via a GPT-4o–assisted vision–language pipeline with human-in-the-loop verification; we release the exact prompts, post-processing rules, and per-field model performance for transparency and reuse.
Beyond flat tables, each frame is encoded as an attribute co-occurrence graph, enabling analyses of semantic structure that go beyond task-specific labels. This representation supports semantic retrieval, scenario mining, and content-aware data selection (e.g., identifying redundancy or rare conditions) without additional model training. We further provide a starter, reproducible baseline with a LoRA-tuned Qwen2-VL-2B with deterministic decoding to calibrate task difficulty via scalar accuracy, micro-averaged F1 for list attributes, and severity MAE/RMSE on a fixed validation split. We publicly release the annotation, and analysis scripts, including graph construction and evaluation scripts, to enable explainable, data-centric workflows for future intelligent vehicles. We provide a demo video as a supplementary material and an anonymized interactive demo (ChatGPT GPT, Drive Scene Analyzer): https://chatgpt.com/g/g-Rpba6Wp9Q-drive-scene-analyzer
II Related Work
Dataset (year) Modality Language supervision Scale Provenance Release type
CARScenes (our work) Single RGB frames Structured scene KB (28 keys; Severity) 5,192 frames 4,692/500 train/val Argoverse1, Cityscapes, KITTI, nuScenes Images + JSONL schema+validator+loader
Talk2Car [RefWorks:deruyttere2019talk2car] (2019) Front camera (nuScenes) Commands with bbox (grounding) 8,349/1,163/ 2,447 cmds nuScenes Text + boxes on nuScenes
nuScenes-QA [RefWorks:qian2024nuscenesqa] (AAAI’24) Multi-modal, multi-frame Programmatic VQA pairs $\sim$460K QAs over $\sim$34K scenes nuScenes Benchmark code + data
NuInstruct [RefWorks:ding2024holistic] (CVPR’24) Multi-view video Instruction–response (17 subtasks) 91K video–QA pairs nuScenes (multi-view) Annotations + loaders
DriveLM-Data [RefWorks:sima2024drivelm] (ECCV’24) Multi-view images/ video Graph-structured QA (P3: perc/pred/plan) n/a nuScenes + CARLA Dataset + eval server
BDD-X [RefWorks:kim2018textual] (2018) Video Action descriptions + explanations 6,970 videos; 77 h $\sim$26K activities BDD Explanations dataset
TABLE I: Related driving VLM datasets compared to CARScenes.
Autonomous-driving datasets. KITTI, Cityscapes, BDD100K, Argoverse, nuScenes, Lyft Level 5, and Waymo Open provide multimodal data with object/pixel-aligned labels for detection, tracking, and mapping. These corpora include limited, coarse scene metadata and are not designed for schema-consistent scene-level semantics. CARScenes complements them by releasing images + JSONL scene annotations for interpretable analysis rather than leaderboard evaluation.
Vision–language for driving. BDD-X, Talk2Car, DriveLM/DriveLM-Data, NuInstruct, and nuScenes-QA pair scenes with free-form text, QA, or graph-QA signals. While powerful, supervision and vocabularies are heterogeneous. CARScenes instead offers a versioned, fixed schema (scene attributes incl. severity) enabling consistent querying and symbolic analysis.
Scene graphs & risk corpora. Visual Genome [RefWorks:krishna2017visual], RefCOCO [RefWorks:chen2025revisiting], and GQA [RefWorks:hudson2019gqa] probe general-image relations/referring or VQA [RefWorks:antol2015vqa], while Scenic [RefWorks:fremont2019scenic], RiskBench [RefWorks:kung2024riskbench], DReyeVE [RefWorks:alletto2016dr], and DeepScenario [RefWorks:lu2023deepscenario] center on simulators, trajectories, or attention—none provide a per-image, driving-specific semantic schema (signals, lane/lane-use, ego context) grounded directly in real images. As a result, they don’t support standardized, per-key evaluation of scene-level attributes or lightweight cross-dataset curation. CARScenes addresses this by offering single-frame, schema-aligned annotations for environment, traffic control, road geometry, actors, and a coarse difficulty/severity proxy, with a consistent evaluation protocol that complements these efforts.
III Dataset Overview
III-A Sources and Scope
CARScenes augments existing autonomous-driving datasets with scene-level annotations and a discrete severity score (1–10). We aggregate 5,192 images from Argoverse1, Cityscapes, KITTI, and nuScenes, chosen for complementary coverage of geography, weather, and urban form. We release annotations and tools only; original imagery remains under the licenses of the source datasets.
III-B Schema and Knowledge Base
Design goals. CARScenes targets scene-level, interpretable supervision that is (i) human-readable and stable across datasets, (ii) VLM-ready for training and evaluation, and (iii) schema-faithful with explicit enumerations and constraints.
III-C 28-Key Category/Sub-Category KB (350+ Attributes)
Group Fields (subset)
Environment TimeOfDay Weather Visibility.General Visibility.SpecificImpairments
Road Geometry & Infrastructure RoadType LaneInformation.NumberOfLanes LaneInformation.SpecialLanes LaneInformation.LaneMarkings IntersectionType
Traffic Control TrafficSigns.Types TrafficSigns.TrafficSignsVisibility TrafficLights.TrafficLightState
Background -Vehicle Behavior Vehicles.TotalNumber Vehicles.VehicleTypes Vehicles.InMotion Vehicles.States
Ego -Vehicle Behavior EgoDirection EgoManeuver
Vulnerable Road Users (VRUs) Pedestrians Cyclists
Sensor States CameraCondition
TABLE II: Overview of the 28-key category/sub-category knowledge base (KB). Full schema includes 350+ leaf attributes.
Each frame also carries a bounded integer Severity on a 1–10 discrete scale.
Field types and naming.
Keys are either scalar categorical (single enum), list-valued (set of attributes), or bounded-integer (Severity). We use dot-delimited names (e.g., TrafficSigns.Types) and canonical enumerations (e.g., Urban, Highway, Night).
Severity definition.
We use severity as an image-only, ordinal difficulty/criticality proxy (1–10) determined from visible cues that plausibly elevate the burden and stakes of scene understanding. The severity score considers (i) road participant complexity (counts, proximity, occlusions), (ii) road/geometry complexity (intersections, merges, lane drops, work zones), (iii) traffic control clarity (visibility/ambiguity of lights, signs, markings), (iv) visibility/conditions (rain, fog, low light, glare, lens artifacts), (v) anomalies/events (stopped vehicles, emergency scenes, debris), and (vi) hazard potential from static cues (e.g., a pedestrian could plausibly emerge from behind a parked vehicle near a crosswalk or on-street parking spaces). Asymmetric consequences are incorporated during labeling: mistakes that would be consequential in context (e.g., treating a red light as green at a signalized junction) increase the assigned severity more than the converse.
Missingness policy.
For list-valued keys, images without that annotation are excluded from list micro-averages; for scalar keys, unannotated fields are omitted from accuracy aggregates. We never conflate “absent” with “unobserved.”
III-D Canonical Record and File Format
All annotations are distributed as JSONL (one record per frame) with exact key names/values validated against a machine-readable schema. A typical record:
"Scene": "Industrial",
"TimeOfDay": "Daytime",
"Weather": "Sunny",
"RoadConditions": "Dry",
"LaneInformation": {
"NumberOfLanes": "Two",
"LaneMarkings": "LaneVisible",
"SpecialLanes": ["NoSpecialLanes"]
"TrafficSigns": {
"TrafficSignsTypes": ["NoTrafficSigns"],
"TrafficSignsVisibility": "SignNotVisible"
"Vehicles": {
"TotalNumber": "MultipleVehicles",
"InMotion": ["False"],
"States": ["Parked"],
"VehicleTypes": ["Pickup Truck", "SUV", "Sedan"]
"Pedestrians": [
"NoPed"
"Directionality": "Two-Way",
"Ego-Vehicle": {
"EgoDirection": "EgoForward",
"EgoManeuver": "EgoMoving"
"Visibility": {
"General": "Good",
"SpecificImpairments": ["NoImpairments"]
"CameraCondition": "Clear",
"Severity": 2
III-E Annotation Pipeline and Quality
Objective. Produce schema-faithful labels that are deterministic, auditable, and suitable for training/evaluating VLMs.
III-F VLM-Assisted Labeling
We prompt a vision–language model with the exact KB and require a single JSON object with a paragraph describing the scene. Decoding is deterministic (temperature 0.0, top-$p$ 1.0, n=1n{=}1). The prompt embeds a schema_version and instructs the model to select only from allowed values.
III-G Post-Processing and Canonicalization
A deterministic validator enforces:
  1. 1.
    Schema checks: type/enumeration validation, dot-delimited keys, canonical case/spelling.
  2. 2.
    Synonym mapping: controlled rewrites from common paraphrases to KB enums.
  3. 3.
    Cross-field constraints: e.g., presence-before-state for TrafficLights.State, mutual exclusions for TimeOfDay, uniqueness for list fields.
  4. 4.
    Normalization: sorted, de-duplicated lists for hash-stable records.
Violations are logged and routed to human review.
III-H Human-in-the-Loop Verification
We use (i) triggered review for any record that fails validation or contains rare/ambiguous combinations and (ii) a stratified random audit by source, weather, and Severity. Reviewers edit via a schema-bound GUI (dropdowns only); edits are logged and become the final labels.
III-I Reliability and Agreement
We quantify label reliability with field-appropriate metrics:
  • Scalar fields: per-key exact-match accuracy and a macro average across scalar keys.
  • List-valued fields: micro-averaged precision/recall/F1 over leaf attributes (denominators exclude images without that field).
  • Severity (1–10): 10-way classification accuracy and regression MAE/RMSE.
IV Dataset Statistics and Structure
This section quantifies the scale and semantic coverage of CARScenes and documents how frames are sampled from source datasets. Our goal is to characterize attribute diversity and severity (1–10) distributions in a way that is directly useful for training and evaluating VLMs.
IV-A Data Sources and Splits
Sources. CARScenes comprises annotated frames from four widely used autonomous-driving datasets: Argoverse1, KITTI, Cityscapes, and nuScenes. These sources jointly cover varied geographies, road types, traffic mixes, and environmental conditions.
Sampling protocol. We uniformly subsample each dataset’s front-facing RGB stream at fixed temporal intervals to reduce near-duplicate frames while preserving transitions between scenarios (e.g., intersection approaches, merges). All selected frames are annotated under the same 28-key knowledge base.
CARScenes includes 5,192 frames sampled from Argoverse1, Cityscapes, KITTI, and nuScenes. We provide a training split of 4,692 images and a fixed validation split of 500 images. Users may re-partition by source/city/condition using provided scripts.
Figure 1: Dataset attribute distribution (scenes and time-of-day).
Figure 2: Severity histogram over all frames.
All multimodal supervision is prepared with utils/qwen_custom_dataset.py using the prompt template defined in configs/qwen_dataset.yaml. The converter pairs each RGB frame with its structured annotation, generating a JSONL record that contains (i) a natural-language analyst summary and (ii) the raw Overall.json label appended verbatim. The resulting training corpus contains 4 6924\,692 samples drawn from four driving benchmarks (Table III) while validation uses a fixed, held-out file at val_dataset_500.jsonl (500 images). Field names appearing in the remainder of this section mirror the keys in Overall.json. No test split is reported here; we encourage future work to reserve additional data for blinded evaluation. We summarize the per-dataset frame counts in Table III and visualize attribute/severity coverage in Figures 12.
TABLE III: Data composition by source. The validation set is disjoint from training and mirrors the geographic/climatic diversity.
Source Train Val
Argoverse1 754 80
Cityscapes 538 57
KITTI 677 72
nuScenes 2 723 291
Total 4 692 500
Figure 1 summarizes the marginal distributions of frequently used keys (e.g., TimeOfDay, Weather, RoadType, Vehicles.TotalNumber, TrafficLights.State). We additionally report the prevalence of list-valued attributes (e.g., Vehicles.Types, Visibility.SpecificImpairments) via micro-averaged support counts.
IV-B Severity (1–10) Distribution
Each frame includes a discrete Severity score in $[1,10]$. Figure 2 shows the histogram over the full dataset and the train/val splits. We also report summary statistics (median, interquartile range) to characterize the concentration of difficulty and to aid stratified sampling.
IV-C Artifacts and Versioning
We release:
  • JSONL annotations with stable key names and canonical enums,
  • a machine-readable schema file and a validator,
  • the full prompt text and post-processing rules,
  • checksums (SHA256) for all files, plus schema_version, prompt_hash, postproc_hash.
IV-D Limitations
Labels are single-frame; certain behaviors (e.g., nuanced maneuvers) can be ambiguous without temporal context. Severity is a coarse, scene-level proxy and does not model downstream risk under a specific ego policy. VLM outputs can reflect pretraining biases; our validation and human review mitigate but do not eliminate such effects.
IV-E Coverage Notes and Imbalance
We highlight coverage and potential imbalance along key axes that commonly affect training and evaluation:
  • TimeOfDay & Weather: distributions reveal daylight dominance in some sources; we recommend stratified sampling for fair comparisons.
  • Traffic control prevalence: TrafficLights.State appears only where lights are present; all list-field denominators exclude images without that annotation.
  • VRU presence: the frequency of Pedestrians/Cyclists varies by source and geography; targeted sampling can improve robustness.
  • Severity spread: mid-range severities are most common; tails (1 and 10) are naturally rarer and useful for stress tests.
IV-F Reproducible Statistics
We provide scripts to regenerate Table III and Figures 12 from the released JSONL files, ensuring that users can recompute all statistics under future schema versions. Outputs include CSV summaries and figure assets in paper/figures/.
IV-G Licensing
CARScenes combines our original materials (annotations, schema, prompt templates, validator, and code) with images from prior datasets (Cityscapes, KITTI, Argoverse v1, nuScenes). We license our annotations/schema/prompts under CC BY 4.0 and our code under MIT; these licenses apply only to our contributions. Images remain under their original licenses and are not re-licensed by us. We do not redistribute Cityscapes or nuScenes images; we provide annotations keyed to official filenames and require users to obtain those images from the original portals under their Terms of Use (nuScenes commercial use requires a separate license from Motional). KITTI and Argoverse v1 images may be redistributed only under their upstream CC BY-NC-SA licenses (KITTI: 3.0; Argoverse v1: 4.0) with attribution; if you include these images, the Non-Commercial + ShareAlike terms flow down to your bundle—otherwise release annotations-only. Consequently, any archive that includes KITTI/Argoverse imagery is non-commercial; for commercial use, exclude those images and pair our annotations with user-downloaded copies from the originals. Please cite the original datasets alongside this work.
V Baseline Vision-Language Model
TABLE IV: VLM MODEL PERFORMANCE (F1) AND INFERENCE TIME (s) AGAINST HUMAN-PROVIDED GROUND TRUTH.
Scene Time Weather Road Lane Signs Vehicles Peds Dir. Ego Vis. Camera Severity Processing Time ChatGPT-4o 0.98 0.99 0.95 0.99 0.95 0.95 0.99 0.98 0.93 0.93 0.99 0.98 0.99 45.63s Qwen2-VL-2B-INS 0.71 0.99 0.74 0.99 0.75 0.76 0.67 0.81 0.88 0.75 0.96 0.99 0.40 44.58s
This section documents the Qwen2-VL-2B-Instruct baseline that accompanies the CARScenes dataset release. Our goal is to provide an immediately reproducible reference model together with the exact training, validation, and evaluation protocol. An abstract of the per-field model performance is given in Table  LABEL:tab:llmvshuman.
V-A Model, Toolkit, and Hyperparameters
We fine-tune Qwen/Qwen2-VL-2B-Instruct with the ms-swift toolkit using a LoRA [RefWorks:hu2022lora] recipe captured in scripts/train_qwen2vl.sh. Table V consolidates the key hyperparameters; the visual encoder remains frozen while LoRA adapters are injected into every language-model projection. Training runs on four 12 GiB 2080Ti GPUs (bf16) for three epochs, resulting in a wall-clock time of $\approx$50.4 hours (1,911 optimizer steps at $\approx$95 s/step, $\approx$202 GPU-hours).
TABLE V: Baseline hyperparameters and schedule. Effective batch size refers to $\texttt{per\_device\_train\_batch\_size}\times\texttt{grad\_accum}$ across four GPUs.
Parameter Value
Base model Qwen/Qwen2-VL-2B-Instruct
Toolkit ms-swift
LoRA rank / $\alpha$ 8 / 32 (decoder $q,k,v,o,\text{up},\text{down},\text{gate}$ proj.)
LoRA dropout / bias 0.05 / none
Visual encoder Frozen (freeze_vit=true)
Max length / max pixels 4 096 tokens / 1 003 520 pixels
[…]
LR schedule Linear warmup (5%) then constant
Grad. clip Disabled
Batch size 1 per GPU, grad. accumulation =2=2 (effective =8=8)
Epochs / steps 3 epochs, 1 911 optimizer steps
Precision bfloat16 (AMP), decoding $T=0.0$, $p=1.0$, $\texttt{max\_new\_tokens}=512$
Validation split Fixed file (500 images; path above)
Seed 42 (ms-swift default)
Selection rule Lowest validation loss (fixed 500-image set)
Run directory output/qwen2vl_lora/v3-20251009-143039
Launch commands.
The helper script exposes all core hyperparameters through environment variables, allowing the baseline to be reproduced with:
CUDA_VISIBLE_DEVICES=0,1,2,3 \
MAX_PIXELS=1003520 \
MAX_LENGTH=4096 \
GRADIENT_ACCUMULATION_STEPS=2 \
scripts/train_qwen2vl.sh
After training, deterministic inference on a folder of frames is available through:
python scripts/infer_qwen2vl_image.py
which reads the prompt template from configs/qwen_dataset.yaml and writes per-image predictions to inference_output/. The evaluation script that produces all numbers in this section can be invoked via:
python scripts/evaluate_validation_set.py \
--val-file output/qwen2vl_lora/v3-20251009-143039/val_dataset_500.jsonl \
--limit None --batch-size 8 --max-new-tokens 512
creating structured_eval_summary.json together with the figures in paper/figures/.
V-B Task Definition and Metrics
The evaluation protocol treats each field in Overall.json as a prediction target. For categorical scalars (e.g., Scene, Directionality) we report top-1 accuracy over the exact class labels provided in the schema. List-valued attributes (e.g., Vehicles.InMotion) are scored with micro-averaged precision, recall, and F1 on the set difference between predicted and ground-truth elements; only images where a field is annotated contribute to the denominator. The discrete severity label (111010) is evaluated both as a classification task (exact match) and as regression using mean absolute error (MAE) and root mean squared error (RMSE). Unless stated otherwise, model selection (Table V) is based on the lowest validation loss measured on the 500-image held-out set.
V-C Validation Results
Table VI lists the scalar-field accuracies on the 500-image validation split (support counts reflect the number of images where the field is annotated; e.g., 162 frames contain traffic lights). Environmental cues (time, road surface, lighting) remain highly reliable (>0.96>0.96), while crowding levels and severity are still the dominant error sources. The corresponding list-valued metrics are shown in Table VII and Figure 3, highlighting strong pedestrian detection ($\mathrm{F1}=0.82$) and visibility impairments ($\mathrm{F1}=0.89$), with moderate performance on lane/sign inventories. Severity regression attains $\mathrm{MAE}=0.72$ and $\mathrm{RMSE}=1.07$, indicating the model continues to overestimate risk by roughly one discrete level.
TABLE VI: Scalar-field accuracy on the validation split.
Field Accuracy Support
Scene 0.81 500
TimeOfDay 0.99 500
Weather 0.82 500
RoadConditions 1.00 500
Directionality 0.83 500
Ego-Vehicle.Direction 0.82 500
Ego-Vehicle.Maneuver 0.79 500
Visibility.General 0.97 500
CameraCondition 0.96 500
TrafficSigns.TrafficSignsVisibility 0.87 500
TrafficSigns.TrafficLightState 0.85 162
Vehicles.TotalNumber 0.79 500
Severity 0.46 500
TABLE VII: List-valued metrics.
Field Precision Recall F1
LaneInformation.SpecialLanes 0.72 0.67 0.70
TrafficSigns.TrafficSignsTypes 0.71 0.66 0.69
Vehicles.VehicleTypes 0.78 0.72 0.75
Vehicles.InMotion 0.81 0.81 0.81
Vehicles.States 0.79 0.72 0.76
Pedestrians 0.83 0.82 0.82
Visibility.SpecificImpairments 0.91 0.87 0.89
Figure 3: Micro-averaged F1 scores for list-valued attributes.
Scalar-field accuracy visualization.
For completeness we include Figure 4, the automatically generated bar chart of per-field accuracy that accompanies the release. These artifacts are committed alongside the code.
Figure 4: Scalar-field accuracy bar chart.
V-D Compute Disclosure and Reproducibility
[…]
  • Runtime: $\approx$50.4 h wall-clock ($\approx$202 GPU-hours; 1,911 steps at $\approx$95 s/step).
  • Software stack: Python 3.10, ms-swift, repository commit, evaluation logs stored in eval_output/native/20251013_121758.
  • Artifacts: Trained LoRA adapters reside in checkpoint-1911; evaluation metrics and confusion matrices are generated via scripts/evaluate_validation_set.py and scripts/plot_.py. The resulting summary (structured_eval_summary.json, is the source of all reported numbers.
Future extensions.
The provided baseline serves as a foundation; we plan to release ablations (e.g., unfreezing the visual encoder, alternative prompt templates) and qualitative panels showing representative successes/failures corresponding to the confusion analysis. We will also host the LoRA checkpoint and a concise reproduction README to facilitate downstream work.
VI Conclusion
We introduced CARScenes, a lightweight, image-grounded semantic dataset for vision–language modeling in autonomous driving that augments 5,192 frames from standard benchmarks with a 28-key schema (350+ attributes) and a discrete Severity label (1–10). Using a deterministic VLM pipeline with human verification, we release the schema, prompts, a validator, and a reproducible baseline intended to calibrate task difficulty (per-key accuracy/F1 and severity regression) rather than to set a leaderboard. CARScenes enables 1) structured-field training/evaluation, compositional generalization over list-valued attributes (e.g., VRUs & night & flashing signal), 2) coverage auditing and curation across time-of-day, weather, and infrastructure, long-tail retrieval and triage via schema queries, cross-dataset transfer under a unified ontology, curriculum/active learning using severity as a difficulty signal, and 3) scenario/test selection by mapping schema constraints to downstream simulation or closed-loop evaluations—all without requiring video or simulators.